A suburban taxi company is considering buying taxis with diesel engines instead of gasoline engines. The cars average 50,000 km a year, with a useful life of 4 years for the taxi with the gas engine and 5 years for the diesel taxi. Other comparative information is as follows:

Diesel                       Gasoline

Vehicle cost                                    $24000                         $19000      

Fuel cost per liter                            $0.68                   $0.72

Mileage, in km/liter                        35                        28

Annual repairs $                    $900                    $700

Annual insurance premiums           $1000                           $1000

End-of-useful-life resale value       $4000                           $6000

Useful life                              5 years                          4 years

Use an annual cash flow analysis to determine the more economical choice if interest is 6%.
Answer #1

Let us take each case one by one

Case 1: Taxi Company buys Diesel Taxi

Intial Vehicle Cost = $24,000

Average distance covered per year= 50,000

Mileage in Km/liter= 35

Therefore liter of diesel required per year = 50,000/35= 1428.57

Fuel Cost per Liter = $0.68

Therefore yearly fuel cost= 1428.57 * 0.68 = $971.43

Annual Repair = $900

Annual Insurance Premium = $1,000

Resale Value after useful life of 5 years = $4,000

So, other than initial vehicle cost, diesel taxi's annual expenses are ($971.43+ $900+ $1000)= $2,871.43

Benefit of selling at year 5 is $4000

Since, interest rate is 6%, discount factor for each year are

Year 0 = 1/(1+0.06)^0 = 1

Year 1= (1.1.06)^1= 0.9434

Year 2= (1/1.06)^2 = 0.8900

Year 3= (1/1.06)^3 = 0.8396

Year 4= (1/1.06)^4 = 0.7921

Year 5= (1/1.06)^5 = 0.7473

So Total Cost for Diesel Taxi is 24,000 + 2871.43 *(0.9434+0.8900+0.8396+0.7921+0.7437) -4,000*0.7473 = $33,106.47

Annual Cashflow will be total cost divided by 5 - (33,106.47)/5= $6,621.29

Case 2: Taxi Company buys Gasoline Taxi

Intial Vehicle Cost = $19,000

Average distance covered per year= 50,000

Mileage in Km/liter= 28

Therefore liter of diesel required per year = 50,000/28= 1785.71

Fuel Cost per Liter = $0.72

Therefore yearly fuel cost= 1785.71 * 0.72 = $1,285.71

Annual Repair = $700

Annual Insurance Premium = $1,000

Resale Value after useful life of 4 years = $6,000

So, other than initial vehicle cost, gasoline taxi's annual expenses are ($1,285.71+ $700+ $1000)= $2,985.71

Benefit of selling at year 4 is $6,000

Discount for Year 1 to Year 4 are shown above. We will use same.

So Total Cost for Gasoline Taxi is 19,000 + 2985.71 *(0.9434+0.8900+0.8396+0.7921) -6,000*0.7921 = $24,593.25

Annual Cashflow will be total cost divided by 4 - (24,593.25)/4= $6,148.31

Since, Annual Cash flow of Gasoline Taxi ($6,148.31) is lesser than that of Desel Taxi ($6,621.29), Gasoline Taxi is more economical choice.

Thanks